X Power 3/2 = 10.41 then x=?
Answer by lwsshak3(11628)

X Power 3/2 = 10.41 then x=?
x^(3/2)=10.41
raise both sides to the 2/3 power
using a calculator
x=10.41^(2/3)≈4.7676